OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = -6.9x + 5.1

OpenStudy (ranga):

The standard form for the equation of a line is Ax + By + C = 0 So first step: Make the right hand side of the equation 0. Then we can go to the next step.

OpenStudy (anonymous):

A. -69x + 10y = 51 B. -69x + 10y = -51 D. 69x + 10y = 51 i know its not c already can u just tell me what it is

OpenStudy (ranga):

I like to teach people how to do it themselves. But if you are just interested in the answers I will just do this one for you but not any more. It is D. 69x + 10y = 51

OpenStudy (anonymous):

thank u I tried to work it out but i couldnt get it.. thanks though